tumi is selling two types of umbrellas.he sells type c at R45 a pair he sells type d at R30 .if he sold 16 pairs in total and had a totl income of R585,how many pairs of sock c did h did Sal sell?

Found 2 solutions by greenestamps, ikleyn:
Answer by greenestamps(13200)



Umbrellas? or pairs of socks?

Tumi? or Sal?

I can't imagine how you would be so careless as to post a question the way you did....

But the meaning is clear: there are two types of items, costing R45 and R30 each; and there are a total of 16 items with a total cost of R585.

The setup for a formal algebraic solution would look something like this:

c = # of items of type
d = # of items of type d

c+d = 16 [the total number of items is 16]
45c+30d = 585 [the total cost is 585]

Multiply the first equation by 30...
30c+30d = 480

... and find the difference between that result and the second equation:

15c = 105
c = 7

ANSWER: c = 7 of type c and 16-7 = 9 of type d

CHECK: 7(45)+9(30) = 315+270 = 585

A quick, informal mental solution -- using exactly the same calculations -- can be obtained using logical reasoning:

If all 16 items were type c, the total cost would be 16(30) = 480
The actual total cost is 585, which is 105 higher
The difference in cost between type c and type d is 15
The number of items of the higher-priced type is 105/15 = 7.


Answer by ikleyn(52790)

.
tumi is selling two types of umbrellas.he sells type c at R45 a pair he sells type d at R30 .
if he sold 16 pairs in total and had a totl income of R585,how many pairs of sock c did h did Sal sell?
~~~~~~~~~~~~~~~~~~~~

Solve using one unknown and one equation.


Let x be the number of items of type "c".
Then the number of items of type "d" is 16-x.


Write the total money equation

    45x + 30*(16-x) = 585  dollars  (revenue).


Divide both sides by 15

    3x + 2(16-x) = 39.


Simplify and find x

    3x + 32 - 2x = 39

    3x - 2x = 39 - 32

       x    =    7.


ANSWER.  7 items of type "c".

Solved.

------------------

To see many other similar problems solved,  look into the lesson
    - Typical problems on buying and selling items
in this site.